A hollow sphere of internal and external radii 3c.m and 4c.m respectively is melted into a cylinder of diameter 37c.m.What is the height of the cylinder ?

Grade:Upto college level

2 Answers

sumanth
10 years ago

given,

       Ri=3

      R2=4

      diameter of the cylinder=37

      radius =18.5

    we know that,volume of the sphere=4/3*3.14*R3

     hence volume of Si=4/3*3.14*3
                                 =36*pie

           volume of Sf=4/3*3.14*33

                               =85.3*pie

    hence volume =Vi-Vf

                        =85.3pie-36*pie

                         =154.9

volume of cylinder=3.14*

 

 

Vinti Singhai
10 years ago
Volume of hollow sphere=Volume of cylinderForm the equation4/3*pi*(4^3-3^3) = pi * 18.5^2*hGet h.Vinti SinghaiaskIITians Faculty,B.E(Honors) in Elec &Comm, Gold Medalist